OpenStudy (anonymous):

laplace ivp x" + 8x' + 15x = 0 , x(0) = 2, x'(0) = -3

OpenStudy (anonymous):

L(x" + 8x' + 15x) = L(0)

OpenStudy (anonymous):

L(0) = 0

hartnn (hartnn):

show what u tried on left side...

OpenStudy (anonymous):

so L(x" + 8x' + 15x) = 0

OpenStudy (anonymous):

so L(x") + 8L(x') + 15L(x) = 0

hartnn (hartnn):

yup. and you know L(x'')

OpenStudy (anonymous):

L(x") = s^2X(s) - sx(0) - x'(0)

hartnn (hartnn):

go on....put initial vales....do L(x') simultaneously.

OpenStudy (anonymous):

what is L(x')

hartnn (hartnn):

L[x'] = s L(x) -x(0)

hartnn (hartnn):

L(x") + 8L(x') + 15L(x) = 0 s^2X(s) - sx(0) - x'(0) + 8( s L(x) -x(0) ) + 15L(x) = 0 put the initial values and isolate L(x)

OpenStudy (anonymous):

i did that but i cant get the answer they have in the back of the book

hartnn (hartnn):

show your work here.

OpenStudy (anonymous):

X(s) = 2s + 13/(s^2 + 8s + 15) where X(s) = L(x)

hartnn (hartnn):

s^2X(s) - 2s +3 + 8 s L(x) - 16 + 15L(x) = 0 thats correct.

hartnn (hartnn):

use partial fraction, denominator can be factorized.

OpenStudy (anonymous):

answer is 1/2(7e^(-3t) - 3e^(-5t))

OpenStudy (anonymous):

okay so X(s) = 2s + 13/(s+5)(s+3)

hartnn (hartnn):

X(s) = (2s + 13)/(s^2 + 8s + 15) = A / (s+5) + B /(s+3)

hartnn (hartnn):

find A,B

OpenStudy (anonymous):

okay so A/(s+5) + B/(s+3) now i multiply A/(s+5) by what

hartnn (hartnn):

u don't know partial fractions ?

hartnn (hartnn):

2s+13 = A(s+3) + B(x+5)

OpenStudy (anonymous):

ist been awhile

hartnn (hartnn):

to get A , put x= -5

hartnn (hartnn):

to get B , put x= -3

hartnn (hartnn):

did u get how i got 2s+13 = A(s+3) + B(x+5)

OpenStudy (anonymous):

2s + 13/(s+5)(s+3) = A/(s+5) + B/(s+3)

OpenStudy (anonymous):

now i cant remember what to do

hartnn (hartnn):

(2s + 13)/(s+5)(s+3) = A/(s+5) + B/(s+3) = [A(s+3)+B(s+5)] / (s+5)(s+3) so 2s+13 = A(s+3) + B(x+5) did u get this ?

OpenStudy (anonymous):

i dont understand

hartnn (hartnn):

which part ? did u get (2s + 13)/(s+5)(s+3) = A/(s+5) + B/(s+3) = [A(s+3)+B(s+5)] / (s+5)(s+3) just cross-,ultiplied.

hartnn (hartnn):

*multiplied

OpenStudy (anonymous):

how you get A(s+3) + B(s+5)

OpenStudy (anonymous):

ohhhh so you multiply both sides by (s+5)(s+3)?

hartnn (hartnn):

i wanted to make the common denominator as (s+5)(s+3) below A there was only s+5 so i multiplied and divided by s+3 A(s+3) / (s+5)(s+3) ok ?

hartnn (hartnn):

yeah.

OpenStudy (anonymous):

it took me awhile but i got it haha

hartnn (hartnn):

good, so can u find A,B?

hartnn (hartnn):

to get A , put x= -5

hartnn (hartnn):

once, u find A,B . just take inverse L.T you'll get the answer.

OpenStudy (anonymous):

wait i have 2s + 13 = (A+B)s + 3A + 5B

hartnn (hartnn):

keep it in this form only 2s+13 = A(s+3) + B(s+5) put s= -5 you'll get A

OpenStudy (anonymous):

WHY DOES s = -5?

hartnn (hartnn):

to find the value of A that equation is satisfied for any value of s so i choose s=-5 so that we eliminate B directly.

OpenStudy (anonymous):

so to find B i plug in s = -3

hartnn (hartnn):

absolutely.

OpenStudy (anonymous):

ive never done it like that before hmm

hartnn (hartnn):

i got to go u find A,B . just take inverse L.T you'll get the answer.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i didnt get it but oh well

hartnn (hartnn):

2s+13 = A(s+3) + B(s+5) put s= -5 3=A(-2) A = -3/2 2s+13 = A(s+3) + B(s+5) put s= -3 7=B(2) B= 7/2 (2s + 13)/(s+5)(s+3) = A/(s+5) + B/(s+3) =(-3/2) / (s+5) + 7/2 (s+3) taking inverse laplace ivp = (-3/2) e^(-5t) + (7/2) e^(-3t) which is the given answer.

hartnn (hartnn):

@amaes